OpenStudy (anonymous):

3x - 4y = 6 6x - 8y = 10 The system shown has _____ solution(s)

OpenStudy (anonymous):

given eqs. are 3x - 4y = 6 and 6x - 8y = 10 for the eqs a1x+b1y = c1 and a2x+b2y = c2 if a1/a2 = b1/b2 <>c1/c2 then the system of equation has no solution..

OpenStudy (anonymous):

thanks